Determine the number of particles in 10.0g of iron
1
Expert's answer
2020-04-10T13:35:34-0400

Solution:

NFe=Na*nFe -the number of particles of iron, here Na=6.02*1023mol-1 - Avogadro constant.

nFe=mFe/AFe=10.0g / 56g*mol-1=0.18mol - the iron amount of substance,

NFe=6.02*1023mol-1*0.18mol=1.08*1023

Answer:

The number of particles in 10.0g of iron is 1.08*1023